Fermi National Laboratory
 

Operation Summary for April 19, 2010 - April 21, 2010

 

Monday April 19

The day shift began with Operations monitoring the Tevatron (TeV) store 7748 (28.3E30), the Recycler stash (48.1E10), and the Antiproton Source stack (42.7E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7748 unintentionally at 8:31 AM. TeV sector D4 cryo system work and a TeV A-E access for various other work began at 9:12 AM. Operators halted MI beam as a precautionary measure from 11:41 AM to 12:53 PM while ComEd performed maintenance work off-site.

During the evening shift, all TeV work was completed by 6 PM. Operators began the TeV turn-on and check-out and put the TeV into a HEP shot setup at 9:16 PM. A problem with a MI quadrupole power supply (I:Q703) load switch halted shot setup at 9:24 PM. Operations called in EE Support personnel to investigate.

Tuesday April 20

The midnight shift began with Operations monitoring the stash (254.6E10) and the Pbar stack (13.2E10), and with NuMI and MiniBooNE off due to load switch work.

During the midnight shift, EE Support personnel had the load switch repaired by 3:25 AM. Operations put the TeV into shot setup and established store 7752 at 4:50 AM with an average initial luminosity (AIL) of 266.2E30.

During the day shift, the Neutron Therapy Facility completed treatment on their last patient at 12:48 PM. Vacuum personnel began work on the MI-11 vacuum system at 1 PM. At 1:29 PM, a TeV RF station (TRF6) tripped off; an RF person investigated. A MI person completed a coalescing check at 1:57 PM.

During the evening shift, an RF person reported at 4:06 PM that he had found and repaired a failed water turbine for the TRF6 power amplifier. A Booster person changed a regulator card for VS4 at 5:13 PM. At 9:13 PM, the High Voltage Coordinator reported a problem with the battery system that supplies power to the DC breaker controls for the Master Substation; a private contractor will investigate and hopefully fix the problem on Wednesday. Operations terminated store 7752 at 9:35 PM. Operations established store 7753 at 11:08 PM with an AIL of 362.1E30; this luminosity was later corrected to 369.1E30.

Wednesday April 21

The midnight shift began with Operations monitoring store 7753 (268.64E30) and the stash (55.9E10), with Pbar stacking (25.6E10), and with NuMI taking beam.

During the midnight shift, Operations reported at 5:29 AM that the I- Source sparked; normal intensity returned. A Linac RF station (LRF2) tripped off; it reset.

The Plans for Wednesday

The plans for today are stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/21/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory